# Personalized Subgraph Federated Learning with Differentiable Auxiliary Projections

### v1

@@ 3.2 Server-Side Federated Aggregation
Thus, after the server performs these personalized aggregations for both $\theta$ and $\bm{a}$, it transmits $(\theta^{(t)}_{k},\bm{a}^{(t)}_{k})$ back to client $G_{k}$ for the $(t+1)$-th communication round starting point. Appendix B shows the pseudo code of FedAux and the complexity analysis.
@@ 3.3 Theoretical Analysis
For notational simplicity, we focus on a single client with $N$ nodes in a given communication round and omit the subscript $k$ and superscript $(t-1)$. The core of our model is to use a learnable auxiliary projection vector APV $\bm{a}$ to capture an optimal node sorting of the local node embeddings and thus serves as a compact summary of the subgraph. However, there is a foundational question that inevitably arises once we replace hard sorting with the smooth kernel aggregator: when the APV $\bm{a}$ is learned via back-propagation, what does it actually learn? Does it encode an arbitrary nuisance direction, or does it converge to a geometrically meaningful axis that faithfully summarises the local subgraph? To answer these questions, we analyze the fidelity of the APV with the following theorem.
@@
where the remainder term obeys $\|\mathcal{R}(\sigma)\|=\mathcal{O}(\sigma^{0})$ as $\sigma\to 0^{+}$. Define $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:\|x\|_{2}=1\}$ as the unit Euclidean sphere embedded in $\mathbb{R}^{d}$, then the gradient descent on $\mathcal{L}$ with unit-norm re-normalization reproduces Oja learning rule (Oja, 1982):
@@
The proof and more discussions are provided in Sec. C.1. It guarantees that, once the kernel aggregator makes $\bm{a}$ differentiable, ordinary back-propagation forces APV to align with the direction along with the node embeddings in that client vary the most. Equivalently, the APV $\bm{a}$ is not an arbitrary trainable knob but a statistically optimal, variance-maximizing summary of local embeddings. Thus, the APV $\bm{a}$ is provably the first principal component of the local embeddings.
In Sec. 3.1, we propose a continuous kernel aggregator to replace the hard sort-then-Conv1D pipeline used in earlier work (Liu et al., 2021). To justify that replacement, the following theorem rigorously shows that the new smooth operator degenerates to the old one in the appropriate limit.
Theorem 3.2 (Sorting limit and equivalence to Conv1D).
Let $z_{i}$ be the kernel-smoothed embeddings, and gather them in score order $\widetilde{\mathbf{Z}}=\left[z_{\pi(1)},\cdots,z_{\pi(N)}\right]\in\mathbb{R}^{N\times d^{\prime}}$. The original sorted embeddings $\widetilde{\mathbf{H}}=[h_{\pi(1)},\cdots,h_{\pi(N)}]$ is defined in Eq. 3. Let $\mathbf{W}\in\mathbb{R}^{B\times d^{\prime}}$ be an arbitrary fixed Conv1D kernel with zero padding, and denote $\mathrm{Conv}_{\mathbf{W}}(\mathcal{X})_{t}=\sum^{B}_{\tau=1}\mathbf{W}_{\tau}\mathcal{X}_{t+\tau-\lceil B/2\rceil}$, for any sequence $\mathcal{X}\in\mathbb{R}^{N\times d^{\prime}}$. Then we have:
@@
The proof is in Sec. C.2. Theorem 3.2 indicates that the kernel aggregation followed by Conv1D converges to hard-sorting followed by conv1D as the bandwidth $\sigma$ tends to 0. Hence, the two architectures have identical expressive power up to an arbitrarily small error for sufficiently small $\sigma$. Although the limit $\sigma\to 0^{+}$ recovers discrete sorting, a larger $\sigma$ performs a soft neighborhood pooling that can act as a learnable regularizer against over-fitting noisy local orderings. In practice, we find $\sigma=1$ to be effective across all datasets.
@@
Let $\Psi^{(t)}_{k}=(\theta^{(t)}_{k},\bm{a}^{(t)}_{k})$ be the local parameters of client $G_{k}$ at the communication round $t$, and $\Psi^{(t)}=\left[\Psi^{(t)}_{1},\cdots,\Psi^{(t)}_{k}\right]$ collects all local parameters. Assuming 1) every local objective is $\mathscr{L}$-smooth: $\forall\Psi_{k},\Psi^{\prime}_{k}:\left\|\nabla\mathcal{L}_{k}(\Psi_{k})-\nabla\mathcal{L}_{k}\left(\Psi^{\prime}_{k}\right)\right\|\leq\mathscr{L}\left\|\Psi_{k}-\Psi^{\prime}_{k}\right\|$; 2) the stochastic gradients are unbiased ($\mathbb{E}\left[g_{k}\right]=\nabla\mathcal{L}_{k}$) and variance-bounded ($\mathbb{E}\left[\left\|g_{k}-\nabla\mathcal{L}_{k}\right\|^{2}\right]\leq\zeta^{2}$), where $g_{k}:=\nabla_{(\Psi_{k})}\mathcal{L}_{k}$ means the local gradients; 3) each local objective satisfies the $\mu$-PL condition (Polyak, 1963); 4) let $\Omega^{(t)}=[w^{(t)}_{kl}]_{k,l}\in\mathbb{R}^{K\times K}$, the spectral gap $\rho:=\sup_{t}\left\|\Omega^{(t)}-\frac{1}{K}\mathbf{11}^{\top}\right\|_{2}<1$. Let each client perform $Q$ local updates per round, and the learning rate $0<\eta\leq\frac{1}{2\mathscr{L}}$. With any initial parameters $\Psi^{(0)}=(\theta^{(0)},\bm{a}^{(0)})$, we have:
@@
The proof is in Sec. C.3. In Theorem 3.3, the first term decays linearly; the second is the classical SGD variance term; the third is the personalization error and vanishes as $\rho\to 0$. Thus FedAux can linearly descend to a neighborhood of the global optimum.
@@ 4 Experiments
Local 81.30$\pm$0.21 79.94$\pm$0.24 80.30$\pm$0.25 69.02$\pm$0.05 67.82$\pm$0.13 65.98$\pm$0.17 84.04$\pm$0.18 82.81$\pm$0.39 82.65$\pm$0.03
FedAvg 74.45$\pm$5.64 69.19$\pm$0.67 69.50$\pm$3.58 71.06$\pm$0.60 63.61$\pm$3.59 64.68$\pm$1.83 79.40$\pm$0.11 82.71$\pm$0.29 80.97$\pm$0.26
FedProx 72.03$\pm$4.56 60.18$\pm$7.04 48.22$\pm$6.81 71.73$\pm$1.11 63.33$\pm$3.25 64.85$\pm$1.35 79.45$\pm$0.25 82.55$\pm$0.24 80.50$\pm$0.25
FedPer 81.68$\pm$0.40 79.35$\pm$0.04 78.01$\pm$0.32 70.41$\pm$0.32 70.53$\pm$0.28 66.64$\pm$0.27 85.80$\pm$0.21 84.20$\pm$0.28 84.72$\pm$0.31
GCFL 81.47$\pm$0.65 78.66$\pm$0.27 79.21$\pm$0.70 70.34$\pm$0.57 69.01$\pm$0.12 66.33$\pm$0.05 85.14$\pm$0.33 84.18$\pm$0.19 83.94$\pm$0.36
FedGNN 81.51$\pm$0.68 70.12$\pm$0.99 70.10$\pm$3.52 69.06$\pm$0.92 55.52$\pm$3.17 52.23$\pm$6.00 79.52$\pm$0.23 83.25$\pm$0.45 81.61$\pm$0.59
FedGTA 71.26$\pm$2.93 68.33$\pm$1.27 69.24$\pm$0.91 69.39$\pm$0.75 67.34$\pm$1.08 65.29$\pm$1.92 78.47$\pm$0.25 82.79$\pm$0.20 81.92$\pm$0.60
FedSage+ 72.97$\pm$5.94 69.05$\pm$1.59 57.97$\pm$12.6 70.74$\pm$0.69 65.63$\pm$3.10 65.46$\pm$0.74 79.57$\pm$0.24 82.62$\pm$0.31 80.82$\pm$0.25
FED-PUB 83.72$\pm$0.18 81.45$\pm$0.12 81.10$\pm$0.64 72.40$\pm$0.26 71.83$\pm$0.61 66.89$\pm$0.14 86.81$\pm$0.12 86.09$\pm$0.17 84.66$\pm$0.54
FedAux 84.57$\pm$0.39 82.05$\pm$0.71 81.60$\pm$0.64 72.99$\pm$0.82 73.16$\pm$0.29 68.10$\pm$0.35 88.10$\pm$0.16 85.43$\pm$0.20 84.87$\pm$0.42
@@
Local 89.22$\pm$0.13 88.91$\pm$0.17 89.52$\pm$0.20 91.67$\pm$0.09 91.80$\pm$0.02 90.47$\pm$0.15 66.76$\pm$0.07 64.92$\pm$0.09 65.06$\pm$0.05
FedAvg 84.88$\pm$1.96 79.54$\pm$0.23 74.79$\pm$0.24 89.89$\pm$0.83 83.15$\pm$3.71 81.35$\pm$1.04 65.54$\pm$0.07 64.44$\pm$0.10 63.24$\pm$0.13
FedProx 85.25$\pm$1.27 83.81$\pm$1.09 73.05$\pm$1.30 90.38$\pm$0.48 80.92$\pm$4.64 82.32$\pm$0.29 65.21$\pm$0.20 64.37$\pm$0.18 63.03$\pm$0.04
FedPer 89.67$\pm$0.34 89.73$\pm$0.04 87.86$\pm$0.43 91.44$\pm$0.37 91.76$\pm$0.23 90.59$\pm$0.06 66.87$\pm$0.05 64.99$\pm$0.18 64.66$\pm$0.11
GCFL 89.07$\pm$0.91 90.03$\pm$0.16 89.08$\pm$0.25 91.99$\pm$0.29 92.06$\pm$0.25 90.79$\pm$0.17 66.80$\pm$0.12 65.09$\pm$0.08 65.08$\pm$0.04
FedGNN 88.08$\pm$0.15 88.18$\pm$0.41 83.16$\pm$0.13 90.25$\pm$0.70 87.12$\pm$2.01 81.00$\pm$4.48 65.47$\pm$0.22 64.21$\pm$0.32 63.80$\pm$0.05
FedSage+ 85.04$\pm$0.61 80.50$\pm$1.30 70.42$\pm$0.85 90.77$\pm$0.44 76.81$\pm$8.24 80.58$\pm$1.15 65.69$\pm$0.09 64.52$\pm$0.14 63.31$\pm$0.20
FedGTA 85.06$\pm$0.82 84.27$\pm$0.71 79.46$\pm$0.28 89.70$\pm$0.67 76.53$\pm$3.21 82.02$\pm$0.78 65.42$\pm$0.09 64.22$\pm$0.08 63.75$\pm$0.18
FED-PUB 90.25$\pm$0.07 89.73$\pm$0.16 88.20$\pm$0.18 93.20$\pm$0.15 92.46$\pm$0.19 90.59$\pm$0.35 67.62$\pm$0.11 66.35$\pm$0.16 63.90$\pm$0.27
FedAux 90.38$\pm$0.08 89.92$\pm$0.15 88.35$\pm$0.96 93.37$\pm$0.26 92.30$\pm$0.29 90.91$\pm$0.60 68.83$\pm$0.15 68.50$\pm$0.27 65.52$\pm$0.10
@@ 4.1 Experimental Setup
Following previous works (Zhang et al., 2021; Baek et al., 2023; Zhang et al., 2024), we construct distributed subgraphs from benchmark datasets by partitioning each original graph into multiple subgraphs corresponding to individual clients. Specifically, we perform experiments on six widely used datasets, including four citation networks (Cora, CiteSeer, Pubmed (Sen et al., 2008), and ogbn-arxiv (Hu et al., 2020)) and two product co-purchase networks (Amazon-Computer and Amazon-Photo (McAuley et al., 2015; Shchur et al., 2018)). We employ METIS (Karypis, 1997) as our default graph partitioning algorithm, which allows explicit specification of the desired number of subgraphs without overlapping. Based on these datasets, we follow the standard experimental setup used in personalized subgraph federated learning literature (Baek et al., 2023; Zhang et al., 2024). Specifically, for dataset splitting, we randomly sample 20%/40%/40% of nodes from each subgraph for training, validation, and testing, respectively. The only exception is the ogbn-arxiv dataset, due to its significantly larger size. For this dataset, we randomly select 5% of the nodes for training, half of the remaining nodes for validation, and the rest for testing. Dataset statistics and implementation details can be found in Appendix D.
@@
FedAux is compared against several representative federated learning (FL) methods, including general FL baselines: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), and FedPer (Arivazhagan et al., 2019); as well as specialized graph-based FL models  [^1] We exclude weakly privacy-protected baselines such as FedGCN (Yao et al., 2023), GraphFL (Wang et al., 2022), and FedStar (Tan et al., 2023) as these methods leak node embeddings, connectivities, or local gradients.: GCFL (Xie et al., 2021), FedGNN (Wu et al., 2021), FedGTA (Li et al., 2023), FedSage+(Zhang et al., 2021), and FED-PUB(Baek et al., 2023). Additionally, we consider a local variant of our model (Local), where FedAux is trained independently at each client without parameter sharing.
@@ 4.2 Main Results
As summarized in Table 1, FedAux is the only algorithm that wins every dataset–client-count combination, indicating that the APV-driven personalization generalizes from small citation graphs (Cora, CiteSeer) to large-scale, high-dimensional graphs (ogbn-arxiv). Specifically, relative to the strongest competitor in each column, FedAux achieves accuracy improvements ranging from 0.2% to 2.4%. The margin over the canonical FedAvg is more pronounced, with an average gain of 4.5%, underscoring that the proposed auxiliary projection mechanism confers benefits well beyond classical weighted averaging. Further, to quantify statistical heterogeneity (i.e., degree of non-IIDness), we adopt $\xi=\textsc{Jsd}+\textsc{Mmd}$ where the Jensen–Shannon Divergence (Jsd) captures label-distribution skew and the Maximum Mean Discrepancy (Mmd) captures disparities in subgraph structure (formal definition in Sec. D.3). Higher values of $\xi$ indicate stronger non-IIDness. As reported in Table 4, $\xi$ rises monotonically for every dataset as the federation enlarges from 5 to 20 clients, confirming that our partition protocol indeed induces progressively harsher heterogeneity. This trend is mirrored in the performance of all methods, whose accuracies decline with larger client counts. Nevertheless, the drop for FedAux is consistently the smallest: on Cora, accuracy falls by only 2.9%, while FedAvg and FedProx lose nearly 7%. To highlight the contrast, we single out the most IID dataset Pubmed and the most non-IID dataset Amazon-Photo in Table 2. The results under these settings show that FedAux remains the top performer in both extremes, indicating that our model is merely tuned for gentle partitions but retains its edge under pronounced non-IID conditions.

### v2

@@ 3.2 Server-Side Federated Aggregation
Thus, after the server performs these personalized aggregations for both $\theta$ and $\bm{a}$, it transmits $(\theta^{(t)}_{k},\bm{a}^{(t)}_{k})$ back to client $G_{k}$ for the $(t+1)$-th communication round starting point. Appendix˜B shows the pseudo code of FedAux.
Complexity Analysis
For the client side of FedAux, the local GNN embedding generation incurs a complexity of $\mathcal{O}(|E_{k}|d^{\prime})$, and the auxiliary projection from embeddings to the APV results in $\mathcal{O}(N_{k}d^{\prime})$. Besides, the kernel-based embedding aggregation over the 1D space induced by the APV has complexity $\mathcal{O}(N_{k}^{2}d^{\prime})$. Consequently, the per-client complexity is $\mathcal{O}(|E_{k}|d^{\prime}+N_{k}^{2}d^{\prime})$. On the server side, computing the client-wise similarity for personalized federated aggregation involves a complexity $\mathcal{O}(K^{2}d^{\prime})$. Therefore, the total complexity of FedAux per communication round is $\mathcal{O}\left((|E_{k}|+N_{k}^{2}+K^{2})d^{\prime}\right)$.
@@ 3.3 Theoretical Analysis
For notational simplicity, we focus on a single client with $N$ nodes in a given communication round and omit the subscript $k$ and superscript (t1)(t-1). The core of our model is to use a learnable auxiliary projection vector APV $\bm{a}$ to capture an optimal node sorting of the local node embeddings and thus serves as a compact summary of the subgraph. However, there is a foundational question that inevitably arises once we replace hard sorting with the smooth kernel aggregator: when the APV $\bm{a}$ is learned via back‑propagation, what does it actually learn? Does it encode an arbitrary nuisance direction, or does it converge to a geometrically meaningful axis that faithfully summarizes the local subgraph? To answer these questions, we analyze the fidelity of the APV with the following theorem.
@@
where the remainder term obeys $\|\mathcal{R}(\sigma)\|=\mathcal{O}(\sigma^{0})$ as $\sigma\to 0^{+}$. Define $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:\|x\|_{2}=1\}$ as the unit Euclidean sphere embedded in $\mathbb{R}^{d}$, then the gradient descent on $\mathcal{L}$ with unit-norm re-normalization reproduces Oja learning rule [22]:
@@
The proof and more discussions are provided in Section˜C.1. It guarantees that, once the kernel aggregator makes $\bm{a}$ differentiable, ordinary back-propagation forces APV to align with the direction along with the node embeddings in that client vary the most. Equivalently, the APV $\bm{a}$ is not an arbitrary trainable knob but a statistically optimal, variance‑maximizing summary of local embeddings. Thus, the APV $\bm{a}$ is provably the first principal component of the local embeddings.
In Section˜3.1, we propose a continuous kernel aggregator to replace the hard sort-then-$\mathrm{Conv1D}$ pipeline used in earlier work [18]. To justify that replacement, the following theorem rigorously shows that the new smooth operator degenerates to the old one in the appropriate limit.
Theorem 3.2 (Sorting limit and equivalence to $\mathrm{Conv1D}$).
Let $z_{i}$ be the kernel-smoothed embeddings, and gather them in score order $\widetilde{\mathbf{Z}}=\left[z_{\pi(1)},\cdots,z_{\pi(N)}\right]\in\mathbb{R}^{N\times d^{\prime}}$. The original sorted embeddings $\widetilde{\mathbf{H}}=[h_{\pi(1)},\cdots,h_{\pi(N)}]$ is defined in Eq.˜3. Let $\mathbf{W}\in\mathbb{R}^{B\times d^{\prime}}$ be an arbitrary fixed $\mathrm{Conv1D}$ kernel with zero padding, and denote $\mathrm{Conv}_{\mathbf{W}}(\mathcal{X})_{t}=\sum^{B}_{\tau=1}\mathbf{W}_{\tau}\mathcal{X}_{t+\tau-\lceil B/2\rceil}$, for any sequence $\mathcal{X}\in\mathbb{R}^{N\times d^{\prime}}$. Then we have:
@@
The proof is in Section˜C.2. Theorem˜3.2 indicates that the kernel aggregation followed by $\mathrm{Conv1D}$ converges to hard-sorting followed by $\mathrm{Conv1D}$ as the bandwidth $\sigma$ tends to 0. Hence, the two architectures have identical expressive power up to an arbitrarily small error for sufficiently small $\sigma$. Although the limit $\sigma\to 0^{+}$ recovers discrete sorting, a larger $\sigma$ performs a soft neighborhood pooling that can act as a learnable regularizer against over‑fitting noisy local orderings. In practice, we find $\sigma=1$ to be effective across all datasets.
@@
Let $\Psi^{(t)}_{k}=(\theta^{(t)}_{k},\bm{a}^{(t)}_{k})$ be the local parameters of client $G_{k}$ at the communication round $t$, and $\Psi^{(t)}=\left[\Psi^{(t)}_{1},\cdots,\Psi^{(t)}_{k}\right$] collects all local parameters. Assuming 1) every local objective is $\mathscr{L}$-smooth: $\forall\Psi_{k},\Psi^{\prime}_{k}:\left\|\nabla\mathcal{L}_{k}(\Psi_{k})-\nabla\mathcal{L}_{k}\left(\Psi^{\prime}_{k}\right)\right\|\leq\mathscr{L}\left\|\Psi_{k}-\Psi^{\prime}_{k}\right\|$; 2) the stochastic gradients are unbiased ($\mathbb{E}\left[g_{k}\right]=\nabla\mathcal{L}_{k}$) and variance-bounded ($\mathbb{E}\left[\left\|g_{k}-\nabla\mathcal{L}_{k}\right\|^{2}\right]\leq\zeta^{2}$), where $g_{k}:=\nabla_{(\Psi_{k})}\mathcal{L}_{k}$ means the local gradients; 3) each local objective satisfies the $\mu$-PL condition [23]; 4) let $\Omega^{(t)}=[w^{(t)}_{kl}]_{k,l}\in\mathbb{R}^{K\times K}$, the spectral gap $\rho:=\sup_{t}\left\|\Omega^{(t)}-\frac{1}{K}\mathbf{11}^{\top}\right\|_{2}<1$. Let each client perform $Q$ local updates per round, and the learning rate $0<\eta\leq\frac{1}{2\mathscr{L}}$. With any initial parameters $\Psi^{(0)}=(\theta^{(0)},\bm{a}^{(0)})$, we have:
@@
The proof is in Section˜C.3. In Theorem˜3.3, the first term decays linearly; the second is the classical SGD variance term; the third is the personalization error and vanishes as $\rho\to 0$. Thus FedAux can linearly descend to a neighborhood of the global optimum.
@@ 4 Experiments
[…]
@@
[…]
@@ 4.1 Experimental Setup
Following previous works [40, 3, 42], we construct distributed subgraphs from benchmark datasets by partitioning each original graph into multiple subgraphs corresponding to individual clients. Specifically, we perform experiments on six widely used datasets, including four citation networks (Cora, CiteSeer, Pubmed [25], and ogbn-arxiv [10]) and two product co-purchase networks (Amazon-Computer and Amazon-Photo [19, 26]). We employ METIS [12] as our default graph partitioning algorithm, which allows explicit specification of the desired number of subgraphs without overlapping. Based on these datasets, we follow the standard experimental setup used in personalized subgraph federated learning literature [3, 42]. Specifically, for dataset splitting, we randomly sample 20%/40%/40% of nodes from each subgraph for training, validation, and testing, respectively. The only exception is the ogbn-arxiv dataset, due to its significantly larger size. For this dataset, we randomly select 5% of the nodes for training, half of the remaining nodes for validation, and the rest for testing. Dataset statistics and implementation details can be found in Appendix˜D.
@@
FedAux is compared against several representative federated learning (FL) methods, including general FL baselines: FedAvg [20], FedProx [14], and FedPer [2]; as well as specialized graph-based FL models  [^1] We exclude weakly privacy-protected baselines such as FedGCN [37], GraphFL [32], and FedStar [30] as these methods leak node embeddings, connectivities, or local gradients.: GCFL [35], FedGNN [34], FedGTA [16], FedSage+[40], and FED-PUB[3]. Additionally, we consider a local variant of our model (Local), where FedAux is trained independently at each client without parameter sharing.
@@ 4.2 Main Results
As summarized in Table˜1, FedAux is the only algorithm that wins every dataset–client-count combination, indicating that the APV-driven personalization generalizes from small citation graphs (Cora, CiteSeer) to large-scale, high-dimensional graphs (ogbn-arxiv). Specifically, relative to the strongest competitor in each column, FedAux achieves accuracy improvements ranging from 0.2% to 2.4%. The margin over the canonical FedAvg is more pronounced, with an average gain of 4.5%, underscoring that the proposed auxiliary projection mechanism confers benefits well beyond classical weighted averaging. Further, to quantify statistical heterogeneity (i.e., degree of non-IIDness), we adopt $\xi=\textsc{Jsd}+\textsc{Mmd}$ where the Jensen–Shannon Divergence (Jsd) captures label-distribution skew and the Maximum Mean Discrepancy (Mmd) captures disparities in subgraph structure (formal definition in Section˜D.3). Higher values of $\xi$ indicate stronger non-IIDness. As reported in Table˜5 of Section˜D.3, $\xi$ rises monotonically for every dataset as the federation enlarges from 5 to 20 clients, confirming that our partition protocol indeed induces progressively harsher heterogeneity. This trend is mirrored in the performance of all methods, whose accuracies decline with larger client counts. Nevertheless, the drop for FedAux is consistently the smallest: on Cora, accuracy falls by only 2.9%, while FedAvg and FedProx lose nearly 7%. To highlight the contrast, we single out the most IID dataset Pubmed and the most non-IID dataset Amazon-Photo in Table˜2. The results under these settings show that FedAux remains the top performer in both extremes, indicating that our model is merely tuned for gentle partitions but retains its edge under pronounced non-IID conditions.
@@
Table 3: Attack AUC of MIA.
Dataset FedGNN FedGTA FedAux Cora 0.56 0.54 0.51 Pubmed 0.58 0.55 0.52 ogbn-arxiv 0.55 0.53 0.49
(a) Cora
(b) Pubmed
(c) Amazon-Computer
Figure 4: Transferability of APV.